# Symbiotic Game and Foundation Models for Cyber Deception Operations in Strategic Cyber Warfare

### v1

@@ 2.2 FM-Enabled GMs
A tactical game is designed to be composable with others. This composability can result in a larger game, leading to a multi-phase and multi-stage operational game that describes a more complex cyber operation [105]. Composability can take on different forms: sequential, expansive, or reinforcing. Sequential composability forms a sequential game where the scenario of one game is followed by that of another. It allows the defender to prepare for forthcoming interactions and make informed defense decisions before the scenario changes, potentially shifting the advantage from the defender to the attacker. Expansive composability involves aggregating two games into one larger game with a more complex structure of action spaces, dynamics, or information structures. It can be viewed as a parallel composition where players engage in two scenarios simultaneously. This enables the defender to confront an attacker with augmented capabilities or multiple attackers coordinating to achieve their goals. Reinforcing composability is a type of feedback composability where the outcome of one game affects the outcome of the other and vice versa. It captures gameplay in which one cyber scenario influences another and eventually returns to itself. When the feedback has a negative impact, the defender faces a vicious cycle where failure leads to further losses in the same scenario.
@@
The development of tactical games and their ability to compose, reason, and learn is not limited to online response and interactions with adversaries. Another advantage lies in the development of game-theoretic digital twins capable of simulating hypothetical scenarios and predicting outcomes previously unencountered. The capabilities of game-theoretic digital twins can inform strategic-level decision making, allowing comprehensive planning by evaluating risks associated with both familiar and unfamiliar scenarios, and the reasoning of an optimal mission-driven strategy. Furthermore, at the tactical level, the game-theoretic digital twin can be also used to create conjectures or predictions of the adversarial behaviors, allowing the defender to gain preparatory advantage over the attacker and improving the tactical decision dominance.
Learning and adaptation occur at multiple levels, forming several adaptation loops. Strategic-level adaptation must coordinate with associated adaptations at the tactical level. A rapid top-down translation between strategic-level mission goals and tactical-level tactics is necessary to enable agile responses to changes in the continually evolving threat landscape in cyberspace. This translation necessitates different levels of knowledge representation. At the strategic level, knowledge and information are encoded as semantic objectives, rules, constraints, and feasible sets. At the tactical level, knowledge and information are represented by acquired spatio-temporal observations encoded in numerical values, time-series data, and binary outcomes. Neurosymbolic learning, which integrates symbolic reasoning with data-driven learning, is essential to facilitate top-down and bottom-up coordination of responses to heterogeneous knowledge and information generated at different levels and from different sources.
@@ 3 Foundation Models and Large Language Models and the Synergetic Roles in Cybersecurity
FMs refer to machine learning models that are trained on broad data, generally using self-supervision at scale, such that they can be adapted to a wide range of downstream tasks. FMs differ from existing machine learning (ML) models in two aspects: scale and scope. Scale: FMs generally feature large-scale neural network models with an astronomical number of parameters to be determined by self-supervised training over comparable data. For example, GPT-3, an FM in natural language processing, has 175 billion parameters and is trained on 45TB of text data. Scope: FM, as a generalist, aims to handle multi-modal data and perform a wide range of tasks without being explicitly trained to do so. It is observed that FMs possess few-show generalizability: a handful of demonstrations from new tasks are sufficient for FMs to adapt. In contrast, ML models are specialists, targeting specific tasks and acting on certain types of data, such as image classification. They typically suffer from poor generalization when transferred to tasks unseen in the training stage. In summary, FMs are large-scale generative models with highly resource-intensive pretraining, capable of learning general features and patterns from diverse data sources. Examples of FMs include large language models (LLM) for natural language processing [72], DALL-E for images[78], MusicGen for music [22], and RT-2 for robotic control [110].
@@
Figure 5: The encoder-decoder structure of the Transformer architecture, adapted from [87]. Each input datapoint, after embedding and positional encoding, is fed to the attention module in the encoder part (the right half), which extracts temporal correlation (attention scores) across the input datapoint sequence. The attention scores are then passed to the decoder (the left half) to generate an output sequence auto-regressively, together with additional attention within the output sequence. The attention mechanism is instrumental in descriptive, predictive, and perspective analytics in cybersecurity.
@@
The few-shot generalization ability of FMs stems from the attention mechanism in the transformer architecture [87] shown in Figure 5, which effectively captures long-range temporal dependencies among inputs. As the building block of FMs, the transformer consists of an encode (on the left half) and a decoder (on the right half). The task of the encoder is to map an input sequence from multi-modal high-dimensional data sources to a sequence of latent representations in a unified vector space, which is then fed to the decoder. Upon receiving the encoder’s output, the decoder generates a new output based on its previous outputs, producing an output sequence in an auto-regressive manner. Mathematically, denote the input sequence by ${\mathbf{x}}=\{x_{1},x_{2},\ldots,x_{n}\}$, which, for example, can be a sentence to be translated or a sequence of sensor readings for robot controls. The transformer is a parameterized auto-regressive probabilistic model $\mathcal{P}_{\theta}$ that consumes the input $\mathbf{x}$ and generates an output sequence $\mathbf{y}=\{y_{1},y_{2},\ldots,y_{m}\}$ using previously generated outputs, i.e., $y_{t}\sim\mathcal{P}_{\theta}(\cdot|y_{t-1},\ldots,y_{1};\mathbf{x})$.
@@
The matrix $QK^{\intercal}$ is divided by $\sqrt{d_{k}}$ to prevent the vanishing gradient problem when applying the soft-max row-wise [87]. The entries of the resulting attention matrix represent the correlation among input tokens. Using the translation example, each entry captures how the word to be translated relates to others in the sentence. Both the encoder and decoder rely on the attention mechanism for correlation extraction in the original sentence $\mathbf{x}$ and the translation $\mathbf{y}$, respectively. Yet, their usage of attention bears a subtle difference. The upper off-diagonal triangle part of the resulting matrix $\text{softmax}({QK^{\intercal}}/{\sqrt{d_{k}}})\in\mathbb{R}^{l\times l}$ is masked with 0s in the decoder part. This causal mask prevents future tokens from influencing the prediction of the current output and is the defining feature of a causal transformer, distinctive from other variations.
To allow the attention mechanism to extract correlation from different representations, FMs often utilize multi-head attention where queries, keys, and values are linearly projected multiple times (one projection corresponds to one head attention) with different learned projections to $d_{k}$, $d_{k}$, and $d_{v}$ dimensions, respectively. The attention mechanism reviewed above acts on the projection in parallel, producing $d_{v}$-dimensional values to be concatenated into a single final value, as shown in Figure 5. The positional encoding and generation of $Q$, $K$, $V$ matrices, together with feed-forward networks and linear layers in the transformer, are trainable components. The parameter size soars up and easily attains the billion level for recent FMs [72], as the embedding dimension, context length, and number of heads increase.
@@ 3.1 Descriptive Analytics
Another emerging game-theoretic security paradigm brought by this descriptive analytics is the natural game representation. Unlike the matrix and other standard game representations in the cybersecurity context, the natural representation features an end-to-end treatment of the actual network systems through FMs and, in particular, large language models (LLM) as the prevalent data in cybersecurity are of text type. LLMs are capable of summarizing evidence-based insights from lengthy cyber threat intelligence (CTI) texts and deriving semantic knowledge from CTI [79], which serves as the game state. Thanks to the attention mechanism, LLMs can discover the correlations between vulnerabilities and attack patterns, mapping Common Vulnerabilities and Exposures (CVE) to CommonWeaknesses Enumeration (CWE) [23] and linking CVEs to MITRE ATT & CK techniques [32]. Such LLM-based mapping naturally specifies the attacker strategy space without undergoing mathematical modeling. The end-to-end security pipeline centered around LLMs and FMs is projected to take an increasing share in cybersecurity task automation.
@@ 3.2 Predictive Analytics
FMs’ predictive power proves to be instrumental in predictive reinforcement learning in the cybersecurity context [35], where the policies are updated online according to the defender’s forecast of the future consequences and anticipation of the attacker’s reactions. Unlike offline or batch reinforcement learning [3], this predictive learning paradigm emphasizes the learning agent’s online adaptability to a subjective forecast of future environment evolution and opponent [58]
@@
Two primary challenges arise from the predictive learning paradigm, to which FMs offer a unified data-driven approach. The first is to adapt the policy on the fly to the forecast with lightweight computation. Existing efforts utilize the multistep lookahead idea whereby the agent predicts multiple rounds of possible interactions (trajectories) in the future [59, 35]. A policy improvement is then carried out by seeking the optimal policy maximizing the cumulative utilities within the lookahead horizon. As pointed out in [59], such a lookahead optimization is more involved than vanilla policy optimization. The adapted policy shall only target a set of plausible forecasts, i.e., those close to the actual trajectory evolution, while discarding all other candidates with marginal probabilities. Consequently, the lookahead adaptation boils down to constrained stochastic optimization, requiring sophisticated machinery for efficient computation.
@@
FMs, as a generative model, naturally fit the role of the predictor in generating future trajectories. Unlike the Monte Carlo simulation that is commonly used in existing works [35, 58], FMs can extract the temporal correlation among past interactions via attention scores and apply them to generate new forecasts without explicitly modeling the environment, leading to high-resolution model-free predictions that are indistinguishable from actual trajectories [52]. The high-accuracy predictions pave the way for online planning algorithms, such as Monte Carlo tree search and rollout methods considered in [59] and [35].
@@
The second challenge pertains to the agent’s subjective perceptions of the environment and the opponent, based on which it makes forecasts. The agent needs to calibrate its subjective conjecture on the environment dynamics [35] and the opponent’s strategy [58] using the online information feedback, e.g., observations. Such a calibration process, performed also online, aims to ensure consistency between the learning agent’s subjective perceptions and the objective multi-agent decision-making [58]: what one observes does not contradict what one believes. To ensure consistency, existing works resort to Bayesian learning approaches, inspired by Bayesian parametric statistics [59, 35, 58]. The proposed Bayesian learning begins with a parametric representation of external uncertainties regarding the environment and the opponent: a set of parameterized models is available to the learning agent, each of which represents a possible system dynamics and opponent strategy. Starting with a prior distribution over these models, the agent continuously updates the posterior upon receiving information feedback, which proves to concentrate asymptotically on the models closest to the actual environment dynamics and opponent response [35, 58].
@@
FMs can revolutionize the calibration process in predictive reinforcement learning. To begin with, it is natural to treat different FMs as parametric models in Bayesian learning, with each pre-trained on different data sources capturing various temporal correlations in system trajectories and opponent sequential actions. Beyond such a straightforward extension, FMs can directly cater to the calibration process in a model-free manner, as investigated in [57]. Without building parametric models beforehand, FMs cut straight to the opponent’s action sequence predictions (forecasts), which are then fed to FMs again to generate the ego agent’s actions conditional on its predictions. The calibration takes place implicitly in the transformer’s auto-regressive operation: the updated forecasts depend on the previous predictions and the information feedback. Recall that the transformer corresponds to a probabilistic model $y_{t}\sim\mathcal{P}_{\theta}(\cdot|y_{t-1},\ldots,y_{1};\mathbf{x})$. In the context of predictive reinforcement learning, $\mathbf{x}$ denotes the past information feedback and the output $y_{t}:=(\widehat{s}_{t:t+K},\widehat{a}_{t:t+K},a_{t})$ includes the the agent’s forecast of future $K$-step system evolution $\widehat{s}_{t:t+K}$, the opponent’s action sequence $\widehat{a}_{t:t+K}$, and the agent’s best response $a_{t}$ against predicted opponent’s moves.
@@
More than a forecaster that generates future trajectories, FMs also assume the roles of the actor (policy improvement) and critic (policy evaluation) in reinforcement learning, leading to an integrated forecaster-actor-critic (FAC) pipeline [58] embodied by the transformer. In contrast to model-based predictive learning [35, 58], FM-powered FAC opens promising avenues to model-free predictive reinforcement learning, exploring a rich class of opponent/system modeling in a data-driven fashion, alleviating the model misspecification in Bayesian learning [35].
@@ 3.3 Prescriptive Analytics
The use of prescriptive analytics, together with descriptive and predictive analytics, extends to the design of sophisticated learning algorithms. One emerging paradigm that can benefit from FMs’ prescriptive capabilities is the non-equilibrium learning [75, 74] and its associated learning defense. Unlike conventional game-theoretic learning, non-equilibrium learning shifts the focus from long-term equilibrium-seeking to transient strategic interaction characterization. Such a paradigm shift is particularly relevant to cybersecurity, where the defender must acquire an advantageous position within a short window before the attack cyber kill chain materializes [60]. As delineated in [75], the three pillars of non-equilibrium learning are the target set (e.g., advantageous positions), the measurement function, and the time window. A learning process is said to achieve non-equilibrium if the measurement of the learning dynamics falls in the target set within the time window. Compared with equilibrium-focused multi-agent learning, non-equilibrium learning emphasizes finite-time approachability to a set of desired outcomes [61]. Thanks to the measurement function, the approachability evaluation in non-equilibrium learning is much more flexible than the utility-driven one in standard game-theoretic learning, which takes into account broader defense objectives than simply utility-maximization, such as privacy and transparency preserving [77, 64].
@@
Naturally, designing learning dynamics from scratch may take a prolonged period, and the defense may completely miss the window of cyber advantage when the defender takes charges before the attack cycle completes. A more viable and effective defense prescription to achieve non-equilibrium is to consider the meta-learning prescription, whereby the FMs first extract common defense strategies or defense response patterns in the offline stage by consuming diverse security data at scale. When deployed online, a handful of online observations on the network system suffices for fast defense policy adaptation, as FMs are decent few-shot learners due to the attention mechanism [72]. The offline pre-trained defense is referred to as the meta-policy, pertaining to a wide variety of security scenarios. As demonstrated in [28, 73], such as meta-learning prescription is scenario-agnostic in the sense that FMs need not examine the exact system configuration. The online adaptation is purely data-driven with marginal computational overhead, leading to effective non-equilibrium defense under system uncertainties.
As we delve into these advanced methodologies, it becomes imperative to reassess traditional equilibrium concepts that underpin our understanding of game outcomes, as motivated by the non-equilibrium concept [75]. The conventional objective has been to compute the equilibrium of the game, where the player’s uncertainties regarding the environment and opponents are internalized as incomplete information [69]. However, the advent of FMs, where agents harness data in diverse ways to formulate strategies, challenges the adequacy of solutions that internalize all uncertainties. Instead, practicality dictates the exploration of solutions corresponding to models that externalize uncertainties.
@@
As such, the pursuit of equilibrium defense is no longer legitimate in cybersecurity, due to the presence of externalities. In this context, the need for novel solution concepts arises. As discussed in the preceding subsection, the notion of consistency [58], inspired by the Berk-Nash [24] and self-confirming equilibrium [26], offers a promising avenue for describing game outcomes in models that explicitly externalize uncertainties, achieving superb performance in reinforcement learning [57] and cybersecurity benchmarks [35]. In addition to consistency, another relevant notion is decision dominance [60], where the dominant party refers to one who can extract information from data and learn about external uncertainties. Whoever completes the information acquisition and analysis cycle faster gains an information advantage over the opponent on the external uncertainties. Unlike classical equilibrium defense resting on the steady state of strategic interactions, decision dominance defense focuses on establishing an information advantage within a transient window and acting decisively before losing the initiative. This shift in perspective acknowledges that not all uncertainties need to be explicitly modeled within the game framework.

### v2

@@ 2.2 FM-Enabled GMs
A tactical game is designed to be composable with others. This composability can result in a larger game, leading to a multi-phase and multi-stage operational game that describes a more complex cyber operation Zhu and Rass (2018). Composability can take on different forms: sequential, expansive, or reinforcing. Sequential composability forms a sequential game where the scenario of one game is followed by that of another. It allows the defender to prepare for forthcoming interactions and make informed defense decisions before the scenario changes, potentially shifting the advantage from the defender to the attacker. Expansive composability involves aggregating two games into one larger game with a more complex structure of action spaces, dynamics, or information structures. It can be viewed as a parallel composition where players engage in two scenarios simultaneously. This enables the defender to confront an attacker with augmented capabilities or multiple attackers coordinating to achieve their goals. Reinforcing composability is a type of feedback composability where the outcome of one game affects the outcome of the other and vice versa. It captures gameplay in which one cyber scenario influences another and eventually returns to itself. When the feedback has a negative impact, the defender faces a vicious cycle where failure leads to further losses in the same scenario.
@@
The development of tactical games and their ability to compose, reason, and learn is not limited to online response and interactions with adversaries. Another advantage lies in the development of game-theoretic digital twins capable of simulating hypothetical scenarios and predicting outcomes previously encountered Li et al. (2024). The capabilities of game-theoretic digital twins can inform strategic-level decision-making, allowing comprehensive planning by evaluating risks associated with both familiar and unfamiliar scenarios and the reasoning of an optimal mission-driven strategy. Furthermore, at the tactical level, the game-theoretic digital twin can also be used to create conjectures or predictions of the adversarial behaviors, allowing the defender to gain a preparatory advantage over the attacker and improving the tactical decision dominance.
Learning and adaptation occur at multiple levels, forming several adaptation loops. Strategic-level adaptation must coordinate with associated adaptations at the tactical level. A rapid top-down translation between strategic-level mission goals and tactical-level tactics is necessary to enable agile responses to changes in the continually evolving threat landscape in cyberspace. This translation necessitates different levels of knowledge representation. At the strategic level, knowledge and information are encoded as semantic objectives, rules, constraints, and feasible sets Peng et al. (2021); Yin et al. (2024). At the tactical level, knowledge and information are represented by acquired spatio-temporal observations encoded in numerical values, time-series data, and binary outcomes. Neurosymbolic learning, which integrates symbolic reasoning with data-driven learning, is essential to facilitate top-down and bottom-up coordination of responses to heterogeneous knowledge and information generated at different levels and from different sources.
@@ 3 Foundation Models and Large Language Models and the Synergetic Roles in Cybersecurity
FMs refer to machine learning models that are trained on broad data, generally using self-supervision at scale, such that they can be adapted to a wide range of downstream tasks. FMs differ from existing machine learning (ML) models in two aspects: scale and scope. Scale: FMs generally feature large-scale neural network models with an astronomical number of parameters to be determined by self-supervised training over comparable data. For example, GPT-3, an FM in natural language processing, has 175 billion parameters and is trained on 45TB of text data. Scope: FM, as a generalist, aims to handle multi-modal data and perform a wide range of tasks without being explicitly trained to do so. It is observed that FMs possess few-show generalizability: a handful of demonstrations from new tasks are sufficient for FMs to adapt. In contrast, ML models are specialists, targeting specific tasks and acting on certain types of data, such as image classification. They typically suffer from poor generalization when transferred to tasks unseen in the training stage. In summary, FMs are large-scale generative models with highly resource-intensive pretraining, capable of learning general features and patterns from diverse data sources. Examples of FMs include large language models (LLM) for natural language processing OpenAI (2023), DALL-E for imagesRamesh et al. (2021), MusicGen for music Copet et al. (2024), and RT-2 for robotic control Zitkovich et al. (2023).
@@
Figure 5: The encoder-decoder structure of the Transformer architecture, adapted from Vaswani et al. (2017). Each input datapoint, after embedding and positional encoding, is fed to the attention module in the encoder part (the right half), which extracts temporal correlation (attention scores) across the input datapoint sequence. The attention scores are then passed to the decoder (the left half) to generate an output sequence auto-regressively, together with additional attention within the output sequence. The attention mechanism is instrumental in descriptive, predictive, and perspective analytics in cybersecurity.
@@
The few-shot generalization ability of FMs stems from the attention mechanism in the transformer architecture Vaswani et al. (2017) shown in Figure 5, which effectively captures long-range temporal dependencies among inputs. As the building block of FMs, the transformer consists of an encode (on the left half) and a decoder (on the right half). The task of the encoder is to map an input sequence from multi-modal high-dimensional data sources to a sequence of latent representations in a unified vector space, which is then fed to the decoder. Upon receiving the encoder’s output, the decoder generates a new output based on its previous outputs, producing an output sequence in an auto-regressive manner. Mathematically, denote the input sequence by ${\mathbf{x}}=\{x_{1},x_{2},\ldots,x_{n}\}$, which, for example, can be a sentence to be translated or a sequence of sensor readings for robot controls. The transformer is a parameterized auto-regressive probabilistic model $\mathcal{P}_{\theta}$ that consumes the input $\mathbf{x}$ and generates an output sequence $\mathbf{y}=\{y_{1},y_{2},\ldots,y_{m}\}$ using previously generated outputs, i.e., $y_{t}\sim\mathcal{P}_{\theta}(\cdot|y_{t-1},\ldots,y_{1};\mathbf{x})$.
@@
The matrix $QK^{\intercal}$ is divided by $\sqrt{d_{k}}$ to prevent the vanishing gradient problem when applying the soft-max row-wise Vaswani et al. (2017). The entries of the resulting attention matrix represent the correlation among input tokens. Using the translation example, each entry captures how the word to be translated relates to others in the sentence. Both the encoder and decoder rely on the attention mechanism for correlation extraction in the original sentence $\mathbf{x}$ and the translation $\mathbf{y}$, respectively. Yet, their usage of attention bears a subtle difference. The upper off-diagonal triangle part of the resulting matrix $\text{softmax}({QK^{\intercal}}/{\sqrt{d_{k}}})\in\mathbb{R}^{l\times l}$ is masked with 0s in the decoder part. This causal mask prevents future tokens from influencing the prediction of the current output and is the defining feature of a causal transformer, distinctive from other variations.
To allow the attention mechanism to extract correlation from different representations, FMs often utilize multi-head attention where queries, keys, and values are linearly projected multiple times (one projection corresponds to one head attention) with different learned projections to $d_{k}$, $d_{k}$, and $d_{v}$ dimensions, respectively. The attention mechanism reviewed above acts on the projection in parallel, producing $d_{v}$-dimensional values to be concatenated into a single final value, as shown in Figure 5. The positional encoding and generation of $Q$, $K$, $V$ matrices, together with feed-forward networks and linear layers in the transformer, are trainable components. The parameter size soars up and easily attains the billion level for recent FMs OpenAI (2023), as the embedding dimension, context length, and number of heads increase.
@@ 3.1 Descriptive Analytics
Another emerging game-theoretic security paradigm brought by this descriptive analytics is the natural game representation. Unlike the matrix and other standard game representations in the cybersecurity context, the natural representation features an end-to-end treatment of the actual network systems through FMs and, in particular, large language models (LLM) as the prevalent data in cybersecurity are of text type. LLMs are capable of summarizing evidence-based insights from lengthy cyber threat intelligence (CTI) texts and deriving semantic knowledge from CTI Ranade et al. (2021), which serves as the game state. Thanks to the attention mechanism, LLMs can discover the correlations between vulnerabilities and attack patterns, mapping Common Vulnerabilities and Exposures (CVE) to CommonWeaknesses Enumeration (CWE) Das et al. (2022) and linking CVEs to MITRE ATT & CK techniques Grigorescu et al. (2022). Such LLM-based mapping naturally specifies the attacker strategy space without undergoing mathematical modeling. The end-to-end security pipeline centered around LLMs and FMs is projected to take an increasing share in cybersecurity task automation.
@@ 3.2 Predictive Analytics
FMs’ predictive power proves to be instrumental in predictive reinforcement learning in the cybersecurity context Hammar et al. (2024), where the policies are updated online according to the defender’s forecast of the future consequences and anticipation of the attacker’s reactions. Unlike offline or batch reinforcement learning Bannon et al. (2020), this predictive learning paradigm emphasizes the learning agent’s online adaptability to a subjective forecast of future environment evolution and opponent Li et al. (2024)
@@
Two primary challenges arise from the predictive learning paradigm, to which FMs offer a unified data-driven approach. The first is to adapt the policy on the fly to the forecast with lightweight computation. Existing efforts utilize the multistep lookahead idea whereby the agent predicts multiple rounds of possible interactions (trajectories) in the future Li et al. (2023); Hammar et al. (2024). A policy improvement is then carried out by seeking the optimal policy maximizing the cumulative utilities within the lookahead horizon. As pointed out in Li et al. (2023), such a lookahead optimization is more involved than vanilla policy optimization. The adapted policy shall only target a set of plausible forecasts, i.e., those close to the actual trajectory evolution, while discarding all other candidates with marginal probabilities. Consequently, the lookahead adaptation boils down to constrained stochastic optimization, requiring sophisticated machinery for efficient computation.
@@
FMs, as a generative model, naturally fit the role of the predictor in generating future trajectories. Unlike the Monte Carlo simulation that is commonly used in existing works Hammar et al. (2024); Li et al. (2024), FMs can extract the temporal correlation among past interactions via attention scores and apply them to generate new forecasts without explicitly modeling the environment, leading to high-resolution model-free predictions that are indistinguishable from actual trajectories Janner et al. (2021); Li et al. (2024). The high-accuracy predictions pave the way for online planning algorithms, such as Monte Carlo tree search and rollout methods considered in Li et al. (2023) and Hammar et al. (2024).
@@
The second challenge pertains to the agent’s subjective perceptions of the environment and the opponent, based on which it makes forecasts. The agent needs to calibrate its subjective conjecture on the environment dynamics Hammar et al. (2024) and the opponent’s strategy Li et al. (2024) using the online information feedback, e.g., observations. Such a calibration process, performed also online, aims to ensure consistency between the learning agent’s subjective perceptions and the objective multi-agent decision-making Li et al. (2024): what one observes does not contradict what one believes. To ensure consistency, existing works resort to Bayesian learning approaches, inspired by Bayesian parametric statistics Li et al. (2023); Hammar et al. (2024); Li et al. (2024). The proposed Bayesian learning begins with a parametric representation of external uncertainties regarding the environment and the opponent: a set of parameterized models is available to the learning agent, each of which represents a possible system dynamics and opponent strategy. Starting with a prior distribution over these models, the agent continuously updates the posterior upon receiving information feedback, which proves to concentrate asymptotically on the models closest to the actual environment dynamics and opponent response Hammar et al. (2024); Li et al. (2024).
@@
FMs can revolutionize the calibration process in predictive reinforcement learning. To begin with, it is natural to treat different FMs as parametric models in Bayesian learning, with each pre-trained on different data sources capturing various temporal correlations in system trajectories and opponent sequential actions. Beyond such a straightforward extension, FMs can directly cater to the calibration process in a model-free manner, as investigated in Li et al. (2023). Without building parametric models beforehand, FMs cut straight to the opponent’s action sequence predictions (forecasts), which are then fed to FMs again to generate the ego agent’s actions conditional on its predictions. The calibration takes place implicitly in the transformer’s auto-regressive operation: the updated forecasts depend on the previous predictions and the information feedback. Recall that the transformer corresponds to a probabilistic model $y_{t}\sim\mathcal{P}_{\theta}(\cdot|y_{t-1},\ldots,y_{1};\mathbf{x})$. In the context of predictive reinforcement learning, $\mathbf{x}$ denotes the past information feedback and the output $y_{t}:=(\widehat{s}_{t:t+K},\widehat{a}_{t:t+K},a_{t})$ includes the the agent’s forecast of future $K$-step system evolution $\widehat{s}_{t:t+K}$, the opponent’s action sequence $\widehat{a}_{t:t+K}$, and the agent’s best response $a_{t}$ against predicted opponent’s moves.
@@
More than a forecaster that generates future trajectories, FMs also assume the roles of the actor (policy improvement) and critic (policy evaluation) in reinforcement learning, leading to an integrated forecaster-actor-critic (FAC) pipeline Li et al. (2024) embodied by the transformer. In contrast to model-based predictive learning Hammar et al. (2024); Li et al. (2024), FM-powered FAC opens promising avenues to model-free predictive reinforcement learning, exploring a rich class of opponent/system modeling in a data-driven fashion, alleviating the model misspecification in Bayesian learning Hammar et al. (2024).
@@ 3.3 Prescriptive Analytics
The use of prescriptive analytics, together with descriptive and predictive analytics, extends to the design of sophisticated learning algorithms. One emerging paradigm that can benefit from FMs’ prescriptive capabilities is the non-equilibrium learning Pan et al. (2023, 2023, 2024) and its associated learning defense. Unlike conventional game-theoretic learning, non-equilibrium learning shifts the focus from long-term equilibrium-seeking to transient strategic interaction characterization. Such a paradigm shift is particularly relevant to cybersecurity, where the defender must acquire an advantageous position within a short window before the attack cyber kill chain materializes Li et al. (2024). As delineated in Pan et al. (2023), the three pillars of non-equilibrium learning are the target set (e.g., advantageous positions), the measurement function, and the time window. A learning process is said to achieve non-equilibrium if the measurement of the learning dynamics falls in the target set within the time window. Compared with equilibrium-focused multi-agent learning, non-equilibrium learning emphasizes finite-time approachability to a set of desired outcomes Li et al. (2021). Thanks to the measurement function, the approachability evaluation in non-equilibrium learning is much more flexible than the utility-driven one in standard game-theoretic learning, which takes into account broader defense objectives than simply utility-maximization, such as privacy and transparency preserving Pawlick et al. (2019); Li and Zhu (2023).
@@
Naturally, designing learning dynamics from scratch may take a prolonged period, and the defense may completely miss the window of cyber advantage when the defender takes charges before the attack cycle completes. A more viable and effective defense prescription to achieve non-equilibrium is to consider the meta-learning prescription, whereby the FMs first extract common defense strategies or defense response patterns in the offline stage by consuming diverse security data at scale. When deployed online, a handful of online observations on the network system suffices for fast defense policy adaptation, as FMs are decent few-shot learners due to the attention mechanism OpenAI (2023). The offline pre-trained defense is referred to as the meta-policy, pertaining to a wide variety of security scenarios. As demonstrated in Ge et al. (2023); Pan et al. (2023), such as meta-learning prescription is scenario-agnostic in the sense that FMs need not examine the exact system configuration. The online adaptation is purely data-driven with marginal computational overhead, leading to effective non-equilibrium defense under system uncertainties.
As we delve into these advanced methodologies, it becomes imperative to reassess traditional equilibrium concepts that underpin our understanding of game outcomes, as motivated by the non-equilibrium concept Pan et al. (2023). The conventional objective has been to compute the equilibrium of the game, where the player’s uncertainties regarding the environment and opponents are internalized as incomplete information Michael et al. (2020). However, the advent of FMs, where agents harness data in diverse ways to formulate strategies, challenges the adequacy of solutions that internalize all uncertainties. Instead, practicality dictates the exploration of solutions corresponding to models that externalize uncertainties.
@@
As such, the pursuit of equilibrium defense is no longer legitimate in cybersecurity, due to the presence of externalities. In this context, the need for novel solution concepts arises. As discussed in the preceding subsection, the notion of consistency Li et al. (2024), inspired by the Berk-Nash Esponda and Pouzo (2016) and self-confirming equilibrium Fudenberg and Levine (1993), offers a promising avenue for describing game outcomes in models that explicitly externalize uncertainties, achieving superb performance in reinforcement learning Li et al. (2023) and cybersecurity benchmarks Hammar et al. (2024). In addition to consistency, another relevant notion is decision dominance Li et al. (2024), where the dominant party refers to one who can extract information from data and learn about external uncertainties. Whoever completes the information acquisition and analysis cycle faster gains an information advantage over the opponent on the external uncertainties. Unlike classical equilibrium defense resting on the steady state of strategic interactions, decision dominance defense focuses on establishing an information advantage within a transient window and acting decisively before losing the initiative. This shift in perspective acknowledges that not all uncertainties need to be explicitly modeled within the game framework.